# OmniRouter: Budget and Performance Controllable Multi-LLM Routing

### v5

@@ A.1 Complete Induction
$L(x,\lambda_{1},\lambda_{2,j},\mu_{i})=\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}\left(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}+\mu_{i}\right)+\lambda_{1}\alpha-\sum_{j=1}^{M}\lambda_{2,j}L_{j}-\sum_{i=1}^{N}\mu_{i}$ (21)
@@
Note that for each $i$, we have the equality constraint $\sum_{j=1}^{M}x_{i,j}=1$. This implies that for each query $i$, exactly one LLM $j$ must be selected. From the stationarity condition, for any fixed $i$, comparing two different indices $j$ and $k$:
$c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}+\mu_{i}=0$ $c_{i,k}-\frac{\lambda_{1}a_{i,k}}{N}+\lambda_{2,k}+\mu_{i}=0$ (26)
@@
$g(\lambda_{1},\lambda_{2})=\min_{x_{i,j}}\{\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j})+\lambda_{1}\alpha-\sum_{j=1}^{M}\lambda_{2,j}L_{j}\}$ (28)
@@
[…]
@@
A.2 Proof
@@
s.t. $\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha$ (38)
$\sum_{i=1}^{N}x_{i,j}\leq L_{j},\quad\forall j\in[M]$ (39)
$\sum_{j=1}^{M}x_{i,j}=1,\quad\forall i\in[N]$ (40)
$x_{i,j}\in\{0,1\},\quad\forall i,j$ (41)
@@
$L(x,\lambda_{1},\lambda_{2},\mu)$ $=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}+\lambda_{1}\left(\alpha-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\right)$ (42)
$\quad+\sum_{j=1}^{M}\lambda_{2,j}\left(\sum_{i=1}^{N}x_{i,j}-L_{j}\right)+\sum_{i=1}^{N}\mu_{i}\left(\sum_{j=1}^{M}x_{i,j}-1\right)$ (43)
@@
Therefore, $x_{i,j^{*}}=1$ and $x_{i,j}=0$ for $j\neq j^{*}$, which completes the proof. ∎
@@
This demonstrates that when $\lambda_{1}^{*}>0$, the quality constraint is satisfied with equality. The dual variable $\lambda_{1}^{*}$ acts as a ”price” for quality: the optimizer finds the solution that exactly meets the quality threshold $\alpha$ while minimizing cost. ∎
Appendix B Dataset Details
B.1 Statistics of the Dataset
The statistics of the OmniRouteEval are in Table 5.
We also set the difficulty for each question, which is determined by the number of models that can answer the question correctly, i.e., Easy: $\{8,9,10\}$ models can answer correctly, Medium: $\{4,5,6,7\}$ models can answer correctly, Hard: $\{0,1,2,3\}$ models can answer correctly. And the proposition is: 78.4%, 15.2%, 6.4%, respectively.
Table 5: Statistics of the data distribution in our dataset.
Data Source # Samples Ratio
MMLU 1000 37.06%
GPQA-Diamond 198 7.33%
Math-500 500 18.53%
GSM8K 1000 37.06%
B.2 Cost Price Map
Table 6: Money Cost Map of LLMs
Model Name Cost per 1M Input Tokens (\$) Cost per 1M Output Tokens (\$) Qwen-2.5-7B-Instruct (Yang et al., 2024) 0.267 0.267 Qwen-2.5-14B-Instruct (Yang et al., 2024) 0.534 0.534 Qwen-2.5-32B-Instruct (Yang et al., 2024) 1.22 1.22 Qwen-2.5-72B-Instruct (Yang et al., 2024) 2.745 2.745 Gemma-2-9B-it (Team et al., 2024) 0.343 0.343 Gemma-2-27B-it (Team et al., 2024) 1.03 1.03 gpt-4o-mini (Achiam et al., 2023) 0.15 0.6 gpt-4o (Achiam et al., 2023) 2.5 10 gemini-1.5-flash (Team et al., 2023) 0.075 0.3 claude-3-5-sonnet [^5] https://www.anthropic.com/news/claude-3-5-sonnet 3 15
B.3 Response Correctness Evaluation
We use the Llama-3.1-70B-Instruct (Dubey et al., 2024) as the LLM judge to evaluate whether a generated response meets the ground truth answers. The prompt is shown as below.
Prompt for using LLM as the judge. Prompt: The ground truth answer is: {gt_answer}. The prediction answer is: extracted_answer. Judge whether the prediction answer is correct or not. You just need to output ’True’ or ’False’.
Appendix C Additional Experiments
C.1 Effects of Key Parameters within OmniRouter’s Predictors.
We investigate the sensitivity of OmniRouter’s routing effectiveness to key design parameters within its prediction components. As shown in Table 7, the number of buckets used for discretizing output length predictions significantly impacts routing performance. A smaller number of buckets (10) yields substantially higher prediction accuracy (45.2% exact match, 80.6% ±$\pm 1$ accuracy), which directly leads to higher routing performance (75.19% success rate) and lower operational costs (\$0.0515). This finding suggests that coarser-grained discretization may be sufficient for effective routing decisions, while excessive granularity can introduce noise. Regarding the retrieval component, Table 8 demonstrates how varying $K$ (the number of historical samples) affects prediction accuracy and routing performance. The system achieves optimal performance at $K=16$, with the highest success rate (75.19%) and lowest cost (\$0.0515). We observe that prediction accuracy remains relatively stable across moderate $K$ values (8-32), with only slight degradation at the extremes. Too few samples ($K=4$) provide insufficient historical context, while too many ($K=64$) likely introduce noise from less relevant queries, both resulting in suboptimal routing decisions. This indicates the retrieval-augmentation requires careful calibration to balance between relevant historical information and potentially misleading outliers.
Appendix D More Related Works
Large Language Models (LLMs) have demonstrated remarkable capabilities across various domains, including mathematical reasoning (Wei et al., 2022; Yao et al., 2024; Besta et al., 2024; Jin et al., 2024), code generation (Wei et al., 2023; Nijkamp et al., 2023; Hui et al., 2024), and recommender systems  Xu et al. (2024a; b; 2023); Mei & Zhang (2023). Many studies (Wei et al., 2022; Yao et al., 2024; Besta et al., 2024; Jin et al., 2024; 2025) have demonstrated that different prompts can have great influences on the model behavior, thus making the prediction of LLM generation length challenging. Early attempts like Magnus (Cheng et al., 2024) employed random forest algorithms but achieved limited accuracy. Subsequent research has explored two main directions of prediction models: encoder-only models for classification (DynamoLLM (Stojkovic et al., 2024), S3 (Jin et al., 2023), TerriInfer (Hu et al., 2024), SSJF (Qiu et al., 2024b), and $\mu$3 (Qiu et al., 2024a)) and decoder-only models for generative prediction like Perception-only (PO) (Zheng et al., 2024). Fu et al. (Fu et al., 2024) reformulated this as a ranking problem and utilized listwise ranking for predictor training. Due to the inherent difficulty in precise output length prediction, several works (Jin et al., 2023; Zheng et al., 2024; Fu et al., 2024) adopted a bucketing approach for approximate estimation. However, these existing studies primarily emphasize computational efficiency while overlooking a critical aspect: model capability - specifically, whether a model has the capability to answer a given query. Therefore, considering both response quality and response cost to construct a general scheduling framework remains a critical challenge.
# Bucket Bucket Acc ($\uparrow$) Capability Acc. ($\uparrow$) \$ Cost ($\downarrow$) Exact Match ±$\pm 1$ 10 0.452 0.806 75.19% 0.0515 20 0.269 0.526 73.22% 0.0725 50 0.162 0.299 73.22% 0.834 100 0.124 0.189 70.74% 0.0962
K Capability Acc. ($\uparrow$) Length Acc ($\uparrow$) SR ($\uparrow$) \$ Cost ($\downarrow$) Exact Match ±$\pm 1$ 4 0.773 0.430 0.784 74.07% 0.0614 8 0.785 0.443 0.786 74.81% 0.0565 16 0.806 0.452 0.806 75.19% 0.0515 32 0.785 0.440 0.776 74.81% 0.0565 64 0.780 0.422 0.764 73.70% 0.0684

### v6

@@ 3.2 Constrained Optimizer
$L(x,\lambda_{1},\lambda_{2,j},\mu_{i})$ (10)
$=\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}\left(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}+\mu_{i}\right)$
$\quad+\lambda_{1}\alpha-\sum_{j=1}^{M}\lambda_{2,j}L_{j}-\sum_{i=1}^{N}\mu_{i}$
@@
Note that for each $i$, we have the constraint: $\sum_{j=1}^{M}x_{i,j}=1$. This implies that for each query $i$, exactly one LLM $j$ must be selected. From the stationarity condition, for any fixed $i$, comparing two different indices $j$ and $k$:
$c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j}+\mu_{i}=0$ (15)
$c_{i,k}-\frac{\lambda_{1}a_{i,k}}{N}+\lambda_{2,k}+\mu_{i}=0$ (16)
@@
$g(\lambda_{1},\lambda_{2})$ (18)
$=\min_{x_{i,j}}\{\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j})+\lambda_{1}\alpha-\sum_{j=1}^{M}\lambda_{2,j}L_{j}\}$
@@
$\max_{\lambda_{1},\lambda_{2}}\quad$ $g(\lambda_{1},\lambda_{2}),$ $\text{s.t.}\quad\lambda_{1}\geq 0,\quad\lambda_{2,j}\geq 0,\quad\forall j\in[M]$ (19)
@@
For $\lambda_{2,j}$, its partial derivative shows the workload violation for each model j. If a model’s assigned queries exceed its capacity, the gradient is positive, increasing $\lambda_{2,j}$, which makes this overloaded model less attractive in subsequent iterations.
4 Evaluation
[…]
4.1 Datasets
We collect 2.7k questions sourced from established knowledge and mathematical reasoning datasets, including MMLU [hendrycks2020measuring], GPQA [rein2023gpqa], MATH [hendrycks2021measuring], and GSM8K [cobbe2021training]. We select 10 different models, including 5 relatively weak models, i.e., Qwen2.5 (7B-Instruct, 14B-Instruct, 32B-Instruct) [yang2024qwen2], Gemma2 (9B-it, 27B-it) [team2024gemma] and 5 relatively strong models, i.e., Qwen2.5-72B-Instruct, GPT 4o-mini, GPT-4o [achiam2023gpt], Gemini-1.5-flash [team2023gemini], Claude-3.5-sonnet  [^1] https://claude.ai/. And we collect the response correctness and token usage of these models to answer these questions using Llama-3.1-70B-Instruct [dubey2024llama] as the evaluation judge [li2024llms]. The Llama-3.1-70B is excluded from the LLM candidate pool to reduce potential biases. The prompt is shown as below.
Prompt for using LLM as the judge. Prompt: The ground truth answer is: {gt_answer}. The prediction answer is: extracted_answer. Judge whether the prediction answer is correct or not. You just need to output ‘True’ or ‘False’.
The statistics of the data we use are in Table 1. We also set the difficulty for each question, which is determined by the number of models that can answer the question correctly, i.e., Easy: $\{8,9,10\}$ models can answer correctly, Medium: $\{4,5,6,7\}$ models can answer correctly, Hard: $\{0,1,2,3\}$ models can answer correctly. And the proposition is: 78.4%, 15.2%, 6.4%, respectively.
Table 1: Statistics of the data distribution in our dataset.
Data Source # Samples Ratio
MMLU 1000 37.06%
GPQA-Diamond 198 7.33%
Math-500 500 18.53%
GSM8K 1000 37.06%
Table 2: Money Cost Map of LLMs
Model Name 1M Input Tokens (\$) 1M Output Tokens (\$) Qwen-2.5-7B-Instruct [yang2024qwen2] 0.267 0.267 Qwen-2.5-14B-Instruct [yang2024qwen2] 0.534 0.534 Qwen-2.5-32B-Instruct [yang2024qwen2] 1.22 1.22 Qwen-2.5-72B-Instruct [yang2024qwen2] 2.745 2.745 Gemma-2-9B-it [team2024gemma] 0.343 0.343 Gemma-2-27B-it [team2024gemma] 1.03 1.03 gpt-4o-mini [achiam2023gpt] 0.15 0.6 gpt-4o [achiam2023gpt] 2.5 10 gemini-1.5-flash [team2023gemini] 0.075 0.3 claude-3-5-sonnet [^3] https://www.anthropic.com/news/claude-3-5-sonnet 3 15
4.2 Setup
We conduct our experiments on an Ubuntu 22.04 machine equipped with 8 RTX A5000 GPUs to serve multiple LLMs using Ollama  [^4] https://ollama.com/. During the implementation, we set the maximum output length to 1024 tokens to prevent the influence of extremely long output sequences, set top $k$ as 16 during retrieval, and $\gamma$ and $\delta$ as 0.5 during aggregation. To calculate money costs, we refer to Litellm’s costmap for price reference. The detailed price map can be found in Appendix B. Our experiments are conducted on the continuous-batching setting [kwon2023efficient]. The serving system processes requests by dynamically forming batches based on the incoming query traffic, providing a more realistic evaluation environment for routing frameworks. In our simulation, we model real-world traffic patterns by randomly adding $n\in\{1,2,3,4\}$ queries to the queue every 0.1 seconds and performing routing decisions at 1-second intervals. If not specifically mentioned, we set the two constraints $\alpha=0.75$ and the concurrent workload constraint $L=4$, which are applied uniformly across all LLMs. Controllability analysis of different values of $\alpha$ and $L$ is presented in Section 4.4. We evaluate OmniRouter against the following routing baselines:
Cost-oriented: We adapt S3 [jin2023s3] and PO [zheng2024response], which are originally designed for latency optimization as cost-oriented baselines. Specifically, S3 employs DistilBERT [sanh2019distilbert] and PO employs Vicuna-7B [chiang2023vicuna] as output token length predictors, respectively.
Performance-oriented: For performance-focused routing, we implement EmbedLLM [zhuang2024embedllm] and RouterDC [chen2024routerdc]. EmbedLLM leverages an encoder-decoder architecture to embed LLM representations, while RouterDC employs contrastive learning to model query-LLM relationships, both aiming to maximize response quality regardless of computational cost.
Cost-performance Coordinated: For baselines which balance both cost and performance objectives, we employ Hybrid-LLM [hybridllm2024], which constructs a probabilistic router, and CARROT [carrot2025], which implements dual predictors using Llama3-8B [dubey2024llama] as specified in its original implementation.
4.3 RQ1: Routing Performance
The comparison of overall routing performance can be observed from Table 3. OmniRouter demonstrates substantial advantages over all baselines across both metrics. Given the theoretical performance bounds (lower bound is 57.41% when the worst LLM for each query is selected and upper bound is 90% when the best LLM for each query is selected), the improvement achieved by OmniRouter (1.3% to 6.48% absolute improvement) is particularly significant, which represents a meaningful advancement within this constrained optimization space. When compared to cost-oriented baselines (S3 and PO), OmniRouter achieves significantly higher success rates (5.74% and 6.48% improvements respectively) while maintaining comparable or even lower costs. Against performance-oriented approaches such as EmbedLLM and RouterDC, our method achieves both higher success rates and substantially lower costs (approximately 41% cost reduction compared to RouterDC). Even when compared to cost-performance coordinated methods (Hybrid-LLM and CARROT), OmniRouter demonstrates superior performance in both dimensions.
Method Capability Acc. Length Bucket Acc. Exact Match $\pm 1$ S3 0.333 0.656 PO 0.325 0.683 EmbedLLM 0.732 RouterDC 0.761 Hybrid-LLM 0.708 CARROT 0.745 0.312 0.672 OmniRouter 0.813 0.452 0.806
Performance of Predictors. We also compare the effectiveness of OmniRouter’s predictor with the baselines we have mentioned. Table 4 demonstrates OmniRouter’s significant advantages in prediction accuracy. Our approach achieves 81.3% capability prediction accuracy, outperforming the strongest baseline (RouterDC) by 5.2%. For length prediction, OmniRouter delivers 45.2% exact match accuracy and 80.6% relaxed ($\pm 1$) accuracy, surpassing all alternatives by over 13.0%.
4.4 RQ2: Controllability Analysis
(a) Cost vs. Performance Constraint
(b) Accuracy vs. Performance Constraint
4.5 RQ3: Ablation Studies
(a) Cost vs. Concurrency Constraint
(b) Accuracy vs. Concurrency Constraint
Predictor Capability Acc ($\uparrow$) Bucket Acc ($\uparrow$) Acc ($\uparrow$) \$ Cost ($\downarrow$) Full 0.813 0.806 75.19% 0.0515 w/o Retrieval 0.751 0.724 71.85% 0.0698 w/o Training 0.728 0.682 73.33% 0.0642
Effects of Key Parameters within OmniRouter’s Predictors. We investigate the sensitivity of OmniRouter’s routing effectiveness to key design parameters within its prediction components. As shown in Table 6, the number of buckets used for discretizing output length predictions significantly impacts routing performance. A smaller number of buckets (10) yields substantially higher prediction accuracy (45.2% exact match, 80.6% $\pm 1$ accuracy), which directly leads to higher routing performance (75.19% success rate) and lower operational costs (\$0.0515). This finding suggests that coarser-grained discretization may be sufficient for effective routing decisions, while excessive granularity can introduce noise. Regarding the retrieval component, Table 7 demonstrates how varying $K$ (the number of historical samples) affects prediction accuracy and routing performance. The system achieves optimal performance at $K=16$, with the highest success rate (75.19%) and lowest cost (\$0.0515). We observe that prediction accuracy remains relatively stable across moderate $K$ values (8-32), with only slight degradation at the extremes. Too few samples ($K=4$) provide insufficient historical context, while too many ($K=64$) likely introduce noise from less relevant queries, both resulting in suboptimal routing decisions. This indicates the retrieval-augmentation requires careful calibration to balance between relevant historical information and potentially misleading outliers.
# Bucket Bucket Acc ($\uparrow$) Capability Acc. ($\uparrow$) \$ Cost ($\downarrow$) Exact Match $\pm 1$ 10 0.452 0.806 75.19% 0.0515 20 0.269 0.526 73.22% 0.0725 50 0.162 0.299 73.22% 0.834 100 0.124 0.189 70.74% 0.0962
K Capability Acc. ($\uparrow$) Length Acc ($\uparrow$) SR ($\uparrow$) \$ Cost ($\downarrow$) Exact Match $\pm 1$ 4 0.773 0.430 0.784 74.07% 0.0614 8 0.785 0.443 0.786 74.81% 0.0565 16 0.806 0.452 0.806 75.19% 0.0515 32 0.785 0.440 0.776 74.81% 0.0565 64 0.780 0.422 0.764 73.70% 0.0684
5 Theoretical Proof
@@
s.t. $\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha\quad\sum_{i=1}^{N}x_{i,j}\leq L_{j},\quad\forall j\in[M]$
$\sum_{j=1}^{M}x_{i,j}=1,\quad\forall i\in[N]\quad x_{i,j}\in\{0,1\},\quad\forall i,j$
@@
$L(x,\lambda_{1},\lambda_{2},\mu)$ (28)
$=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}+\lambda_{1}\left(\alpha-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\right)$
$\quad+\sum_{j=1}^{M}\lambda_{2,j}\left(\sum_{i=1}^{N}x_{i,j}-L_{j}\right)+\sum_{i=1}^{N}\mu_{i}\left(\sum_{j=1}^{M}x_{i,j}-1\right)$
@@
Therefore, $x_{i,j^{*}}=1$ and $x_{i,j}=0$ for $j\neq j^{*}$, which completes the proof.
@@
This demonstrates that when $\lambda_{1}^{*}>0$, the quality constraint is satisfied with equality. The dual variable $\lambda_{1}^{*}$ acts as a ”price” for quality: the optimizer finds the solution that exactly meets the quality threshold $\alpha$ while minimizing cost.
6 Related Work
LLM Routing. To optimize performance and inference cost in multi-LLM serving systems, researchers have developed increasingly sophisticated routing approaches. RouteLLM [ong2024routellm] and CARROT [carrot2025] propose training small LLMs as routers to balance cost and performance. EmbedLLM [zhuang2024embedllm] introduces a specialized encoder-decoder network for embedding LLM representations. HybridLLM [hybridllm2024] addresses data imbalance issues by proposing a probabilistic router to better represent different model capabilities across query types. C2MAB-V [dai2024cost] treats routing as a contextual multi-armed bandit problem, enabling exploration-exploitation trade-offs in model selection. GraphRouter [feng2024graphrouter] leverages graph learning to jointly modeling the query-model, query-query, and model-model relationship for building routers. While these approaches have made substantial progress, they predominantly operate as greedy decision-makers, optimizing for individual queries without considering global system constraints or the implications for subsequent requests. DeepSieve [guo2025deepsieve] utilizes LLM-as-a-Knowledge-Router to enable effective reasoning across multiple heterogeneous knowledge sources, enhancing retrieval-augmented generation (RAG) systems by dynamically routing sub-queries to the most appropriate source.
LLM Generation Length Prediction. Predicting LLM generation length is crucial for optimizing computational resources. Early attempts like Magnus [cheng2024enabling] employed random forest algorithms but achieved limited accuracy. Subsequent research has explored two main directions of prediction models: encoder-only models for classification (DynamoLLM [stojkovic2024dynamollm], S3 [jin2023s3], TerriInfer [hu2024inference], SSJF [aiops2024qiu], and $\mu$3 [qiu2024power]) and decoder-only models for generative prediction like Perception-only (PO) [zheng2024response]. LTR [fu2024efficient] reformulated this as a ranking problem and utilized listwise ranking for predictor training.
7 Conclusion